OpenStudy (tad1):

Find the second derivative of xy +y^2 = 1. The first derivative is -y(x + 2y)^(-1). The text book answer is 2(x + 2y)^-3, but I've been unable to get that answer.

OpenStudy (anonymous):

y+xy'+2yy'=0 y'=-y/(x+2y)

OpenStudy (amistre64):

Dx(xy +y^2 = 1) Dx(xy) +Dx(y^2) = Dx(1) (x'y + xy') +(2y y') = (0) x'y + y'(x +2y) = 0 y'(x +2y) = -x'y y' = -x'y/(x +2y)

OpenStudy (tad1):

Thanks, but I have the first derivative. I need the second.

OpenStudy (amistre64):

since dx/dx = 1 .. x'=1 Dx[y' = -y/(x +2y)] Dx[y'] = Dx[-y (x +2y)^-1] y'' = -y'(x+2y^-1) + -y -(x+2y)^-2 Dx(x+2y)

OpenStudy (amistre64):

Dx(x+2y) Dx(x) + Dx(2y) x' + 2 y' 1 + 2y' \[y'' = -y'(x+2y^{-1}) + y(x+2y)^{-2} (1+2y')\] \[y'' = \frac{-y'}{(x+2y)^{-1}} + \frac{y (1+2y')}{(x+2y)^{-2}}\] \[y'' = \frac{-y'}{(x+2y)^{-1}} + \frac{y+2yy'}{(x+2y)^{-2}}\]

OpenStudy (amistre64):

get like denoms and it should simplify to the text

OpenStudy (amistre64):

but I think I see a slight error in my math to text skills lol

OpenStudy (anonymous):

denominator (x+2y)?

OpenStudy (anonymous):

also(x+2y)2

OpenStudy (amistre64):

you know; Newton was wrong alot too lol

OpenStudy (amistre64):

the 1st derive is good; the second needs som etender lovin care applied to finesse it to the proper look

OpenStudy (anonymous):

the prob is just with denominator

OpenStudy (anonymous):

(x+2y)^2 is the lcm

OpenStudy (amistre64):

\[Dx[y' = -\frac{y}{(x +2y)}]\] \[y'' = -\frac{(x+2y)y' - y(1+2y')}{(x +2y)^2}\]

OpenStudy (anonymous):

y''=(-xy'+y)/(x+2y)^2

OpenStudy (anonymous):

y''=xy/(x+2y)+y ------------ (x+2y)^2

OpenStudy (anonymous):

even then it doesn't match the solution:S all becoz of typing :)

OpenStudy (amistre64):

\[y'' = -\frac{y'x +2yy' -y-2yy'}{(x +2y)^2}\] \[y'' = -\frac{y'x -y}{(x +2y)^2}\] \[y'' = -\frac{(-\frac{y}{(x +2y)})x -y}{(x +2y)^2}\]

OpenStudy (anonymous):

=2xy+2y^2/(x+2y)^3

OpenStudy (amistre64):

\[y'' = -\frac{-\frac{yx}{(x +2y)} -\frac{y(x+2y)}{(x+2y)}}{(x +2y)^2}\] \[y'' = -\frac{-yx - y(x+2y)} {(x+2y)(x +2y)^2}\] \[y'' = -\frac{-yx - yx-2y^2} {(x +2y)^3}\]

OpenStudy (anonymous):

2(xy+y^2)/(x+2y)^3=2(x+2y)^-3

OpenStudy (anonymous):

yayyy...got it :)

OpenStudy (anonymous):

xy+y^2=1

OpenStudy (amistre64):

\[y'' = -\frac{-2yx-2y^2} {(x +2y)^3}\] \[y'' = \frac{2y(x+y)} {(x +2y)^3}\]

OpenStudy (anonymous):

newton in wrong direction again ;)

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

taking 2 outside

OpenStudy (anonymous):

go on, m so bad in latex

OpenStudy (tad1):

Here's one of my attempts y'' = [y(1 + 2 y' - (x +2y)]/(x+2y)^2 y + 2y(-y/(x+2y) - (x+2y)/ (x +2y)^2 = y (x+2y) -2y^2 - (x+2y)^2]/(x +2y)^3 =then I'm having trouble

OpenStudy (tad1):

Thanks to both of you for hanging in there.